Tshego won R165 000 and decided to deposit 65% of this amount in an account earning 8,25% interest per year, compounded every four months. The accumulated amount after five years is


[1] R248 203,55

. [2] R151 490,63.

[3] R247 862,83.

[4] R161 110,84.

[5] R161 332,31


1
Expert's answer
2021-05-24T01:10:03-0400

Solution:

Given, P= R 165,000

Deposit amount = 65% of P

r = 8.25% ,n = 4, t = 5

PV "=65\% \times R\ 165,000"

"\Rightarrow PV =R\ 107250"

"A=p(1+\frac{r}{n})^{nt}"

"\Rightarrow A=107250×(1+\frac{8.25\%}{4})^{4×5}"

"\Rightarrow A=R\ 161332.31"

Hence option (5) is correct.